OpenStudy (anonymous):

Algebra. x * y = 5166, x+y 145. This I know I could do with simply guess and check, but what is the proper algebraic way of doing it? Thanks a lot!

OpenStudy (wolf1728):

A) x * y = 5,166 B) x + y = 145 y = 5,166/x x + 5,166/x = 145 x² + 5,166,= 145x x² -145x + 5,166,= 0 Using quadratic formula to solve: x = [--145 +-sqroot(21,025 - 4*1*5,166)] / 2*1 x = 145 +-sqroot(21,025 -20,664) / 2 x = 145 +-sqroot(21,025 -20,664) / 2 x = 145 +-sqroot(361) / 2 x = 145 +- 19 / 2 x1 = 164 / 2 = 81 x2= 145 -19 / 2 x2 = 126 / 2 x2 = 63

OpenStudy (anonymous):